# Peeking with PEAK: Sequential, Nonparametric Composite Hypothesis Tests for Means of Multiple Data Streams

### v2

@@ 5.3 Discussion of Results
6 Conclusions and Future Directions.
In our work, we propose a novel nonparametric sequential test for composite hypotheses over means of multiple data streams. In the single arm case, our approach provides a robust option among existing sequential tests and anytime valid confidence sequences. In the multi-arm case, our test pools evidence across arms to avoid union bounds, and empirically outperforms other sequential tests both (i) as a stopping criteria for common pure-exploration bandit problems, such as threshold or best-arm identification, and (ii) in terms of runtime compared leading approaches for sequential testing. Importantly, our approach is compatible with various existing convex optimization methods, avoiding the need to perform grid-search style methods as subroutine while still maintaining the performance of other testing-as-betting sequential tests. Future directions for this work include (1) a general approach for arbitrary convex hypothesis regions, which include projected gradient descent methods, (2) developing a sampling scheme tailored towards our test for specific objectives, and (3) generalizing our approach to settings involving non-stationary arms/streams.
Broader Impact
@@ A.1 Proof of Theorem 1
Note that $f(\epsilon,m)=\mathbb{E}[\log((1+\frac{\mu-m_{)}(X_{i}-m)}{c}]>0$ at $\epsilon=0$ (proven above). Because $f(\epsilon,m)$ is continuous with respect to $\epsilon$, we know there exists an $\epsilon(m)$ such that $f(\epsilon(m),m)=\mathbb{E}[\log((1+\frac{(\mu-m)(X_{i}-m)}{c}]/2>0$. By the Kolmogorov Strong Law of Large Numbers, we know there exists a $t(m)\in\mathbb{N}$, $t(m)<\infty$ such that $\hat{\mu}_{t}\in[\mu-\epsilon(m),\mu+\epsilon(m)]$ for all $t\geq t(m)$, for any sample path $\omega$, where $\omega\in\Omega$ s.t. $\mathbb{P}(\Omega)=1$. The log capital process $l_{t}(m)$, for $t>t(m)$, takes the following form:
@@
Summation $(a)$ is guaranteed to be finite by $t(m)<\infty$ for all sample paths, and we will now show that summation $(b)$ diverges to $\infty$ for all $m_{a}\neq\mu$. We can rewrite summation $(b)$ as the following:
@@
$(b)\geq\sum_{i=t(m)}^{t}\mathbf{1}[X_{i}\geq m_{a}]\log(1+\frac{(\mu-\epsilon(m_{a})-m_{a})(X_{i}-m_{a})}{c})+\sum_{i=t(m)}^{t}\mathbf{1}[X_{i}<m_{a}]\log(1+\frac{(\mu+\epsilon(m_{a})-m_{a})(X_{i}-m_{a})}{c}).$
@@
By the SLLN and by the fact that the LHS converges to $f(\epsilon(m),m)/2>0$, there exists a finite time $t^{\prime}(m)\in\mathbb{N}$ such that for all $t\geq t^{\prime}(m)$, the following holds:
@@
A.4 Proof of Lemma 1
Proof of Lemma 1.
We prove this for the case where $\mu>m$; the other case is a symmetric proof. By the SLLN, for all $\epsilon>1$ and for all sample paths $\omega\in\Omega$ (where $\mathbb{P}(\Omega)=1$), there exists a $t(\epsilon)\in\mathbb{N}$, $t(\epsilon)<\infty$, s.t. for all $t\geq t(\epsilon)$,
$\hat{\mu}_{t-1}\geq\mu-(1-1/\epsilon)(\mu-m).$
Pluggin this into our expression, for $t\geq t(\epsilon)$, we obtain that:
$\mathbb{E}_{P}[l_{t}(m)-l_{t-1}(m)]$ $=\mathbb{E}_{P}[\log(1+\frac{(\hat{\mu}_{t-1}-m)(X-m)}{c})]$
$=\mathbb{E}_{P}[\log(1+\frac{\frac{(\hat{\mu}_{t-1}-m)}{(\mu-(1-1/\epsilon)(\mu-m)-m)}*(\mu-(1-1/\epsilon)(\mu-m)-m)(X-m)}{c})]$
$=\mathbb{E}_{P}[\log(1+\frac{(\mu-m)(X-m)}{\epsilon\frac{(\mu-(1-1/\epsilon)(\mu-m)-m)}{(\hat{\mu}_{t-1}-m)}c})]$
$\underbrace{\geq}_{=(i)}\mathbb{E}_{P}[\log(1+\frac{(\mu-m)(X-m)}{\epsilon c})].$
where inequality $(i)$ holds from the fact that if $\hat{\mu}_{t-1}\geq\mu-(1-1/\epsilon)(\mu-m)$, $\mu-m>0$, then
$\frac{(\mu-(1-1/\epsilon)(\mu-m)-m)}{(\hat{\mu}_{t-1}-m)}\leq 1,$
and by Lemma 2, smaller values in the denominator term increase the value of the expectation in the preceding line. For the $\mu<m$ case, we replace the $t(\epsilon)$ condition with $\hat{\mu}_{t-1}\leq\mu+(1-1/\epsilon)(\mu-m)$ to obtain the identical result.
A.5 Proof of Lemma 2
@@
Taking the difference, we obtain the desired result:
@@ A.7 Proof of Theorem 3
Thus, by Ville’s inequality 6, we obtain that the crossing probability $\mathbb{P}(\exists t\in\mathbb{N}:T_{t}(m)=\mathbf{1}[E_{t}(m)\geq 1/\alpha]=1)\leq 1/\alpha$, guaranteeing our desired Type I error coverage. ∎
A.8 Proof of Theorem 4
@@ A.10 Proof of Proposition 3
We provide additional details on confidence sequences used in the single arm experiments in Section 5. All experiments, including runtimes, are given in
@@
The termination condition for a single arm $a$ (i.e. below/above threshold $\epsilon$) is determined by the following anytime valid confidence interval in (Kano et al.,, 2018), which we refer to as $C_{t}^{\text{base}}$:
@@
Runtime Testing.

### v3

@@ 5.1 Synthetic Experiments
5.2 Case Study with Mobile Health Data
We use the HeartSteps dataset (Liao et al.,, 2020) as a case study for our approach. The HeartSteps dataset was collected using a data-adaptive sampling scheme during a mobile health intervention with the HeartSteps app and contains user-level information on 37 users: (1) $A_{i,t}$ a binary intervention received at time $t$ for user $i$, and (2) $X_{i,t}$, the number of user $i$’s steps within a short interval after the intervention. In our case study, we investigate the value of $\tau_{i}$, the first time in the trial in which we can determine whether $\mu_{i}(1)\geq\mu_{i}(0)$ (or vice versa) with 70% confidence. We make the following assumptions for the DGP: (1) each user $i$ has an associated mean vector $\mu_{i}$, where for all $t$, $\mathbb{E}[X_{i,t}|\mathcal{F}_{t-1},A_{i,t}=a]=\mu_{i}(a)$, and (2) the number of steps that can be taken are bounded above for all users. These two assumptions are sufficient for the type-I error guarantees of PEAK. We apply the BAI termination condition for $W=2$ for each $i$ in the dataset using PEAK, Sub-B, and Hedged, the three most empirically promising testing schemes from Section 5.1, to obtain $\tau_{i}$ (if it occurs before the horizon of the data) for each user $i$.
In our experiments, we observe that Sub-B fails to conclude that the intervention is either helpful or harmful for any user $i$ throughout the entire horizon of the trial. In contrast, both testing-by-betting methods are able to conclude that a treatment is helpful/harmful before the end of the trial, to varying degrees. Hedged concludes that the digital intervention provides no benefit and may be harmful (i.e., $\mu_{i}(1)\leq\mu_{i}(0)$) for User 22 using just half of the samples collected over the trajectory ($\tau_{22}=4932$ over a horizon of 9072). For all other patients, Hedged fails to conclude that the intervention is either helpful or harmful. PEAK outperforms both approaches: it provides stopping times $\tau_{i}$ for 8 different patients before the end of the recorded data, identifying 8 different patients who all benefit from the treatment (i.e., $\mu_{i}(1)\geq\mu_{i}(0)$). If the analyst were to stop trials for the 8 users at their respective stopping times, they would have saved 18,563 total samples, reducing the total number of samples by roughly 10%. This simple case study demonstrates PEAK’s potential to reduce experimental costs and provide resources for additional patient enrollment.
6 Conclusions and Future Directions
In our work, we propose a novel nonparametric sequential test for composite hypotheses over means of multiple data streams. In the single-arm case, our approach provides a robust option among existing sequential tests and anytime valid confidence sequences. In the multi-arm case, our test pools evidence across arms to avoid union bounds, and empirically outperforms other sequential tests both (i) as a stopping criterion for common pure-exploration bandit problems, such as threshold or best-arm identification, and (ii) in terms of runtime compared with leading approaches for sequential testing. Importantly, our approach is compatible with various existing convex optimization software, avoiding the need for grid-search style methods while maintaining the performance of other testing-by-betting methods.
Despite improved computational tractability, PEAK requires the use of solvers to test hypotheses; existing parametric $e$-processes have minimal computational overhead relative to any testing-by-betting approach, including PEAK. Our theoretical/empirical results suggest that PEAK is most appropriate for problems where (1) strong parametric assumptions are unrealistic, and (2) one wishes to preserve power against both high-variance and low-variance data streams.
Acknowledgements
Brian Cho was supported by the NDSEG fellowship.
Impact Statement
@@ A.1 Proof of Theorem 1
Note that $f(\epsilon,m)=\mathbb{E}[\log((1+\frac{\mu-m_{)}(X_{i}-m)}{c}]>0$ at $\epsilon=0$ (proven above). Because $f(\epsilon,m)$ is continuous with respect to $\epsilon$, we know there exists an $\epsilon(m)$ such that $f(\epsilon(m),m)=\mathbb{E}[\log((1+\frac{(\mu-m)(X_{i}-m)}{c}]/2>0$. By the Kolmogorov Strong Law of Large Numbers, we know $t(m,\omega)\in\mathbb{N}$, $t(m,\omega)<\infty$ such that $\hat{\mu}_{t}\in[\mu-\epsilon(m),\mu+\epsilon(m)]$ for all $t\geq t(m,\omega)$ almost surely, where $\omega\in\Omega$ s.t. $\mathbb{P}(\Omega)=1$. We denote $t(m,\omega)$ as $t(m)$ to keep notational clutter minimal. The log capital process $l_{t}(m)$, for $t>t(m)$, takes the following form:
@@
Summation $(a)$ is guaranteed to be finite almost surely by $t(m)<\infty$, and we will now show that summation $(b)$ diverges to $\infty$ for all $m_{a}\neq\mu$. We can rewrite summation $(b)$ as the following:
@@
$(b)\geq$ $\sum_{i=t(m)}^{t}\mathbf{1}[X_{i}\geq m_{a}]\log(1+\frac{(\mu-\epsilon(m_{a})-m_{a})(X_{i}-m_{a})}{c})$
$+\sum_{i=t(m)}^{t}\mathbf{1}[X_{i}<m_{a}]\log(1+\frac{(\mu+\epsilon(m_{a})-m_{a})(X_{i}-m_{a})}{c}).$
@@
By the SLLN and by the fact that the LHS converges to $f(\epsilon(m),m)/2>0$, there exists a finite time $t^{\prime}(m,\omega)\in\mathbb{N}$ almost surely for each sample path $\omega\in\Omega$ such that for all $t\geq t^{\prime}(m,\omega)=t^{\prime}(m)$, the following holds:
@@
A.4 Proof of Lemmas 1 and and 3
@@
Taking the ratio, we obtain the desired result:
@@ A.6 Proof of Theorem 3
Thus, by Ville’s inequality 5, we obtain that the crossing probability $\mathbb{P}(\exists t\in\mathbb{N}:T_{t}(m)=\mathbf{1}[E_{t}(m)\geq 1/\alpha]=1)\leq 1/\alpha$, guaranteeing our desired Type I error coverage. ∎
A.7 Proof of Theorem 4
@@ A.9 Proof of Proposition 3
In this section, we provide additional details on the baseline methods of comparison, the data-adaptive sampling schemes used by in Section 5.2, additional details regarding Table 3, and additional simulations using non-standard underlying distributions for each arm.
B.1 Additional Details on Baseline Confidence Sequences / Anytime Valid Tests.
We provide additional details on confidence sequences used in the single arm experiments in Section 5.
@@
The termination condition for a single arm $a$ (i.e. below/above threshold $\epsilon$) is determined by the following anytime valid confidence interval in (Kano et al.,, 2019), which we refer to as $C_{t}^{\text{base}}$:
@@
B.2 Additional Details For Sampling Schemes in Multi-Arm Case.
We provide the exact sampling schemes used for Section 5’s experiments regarding threshold identification (THR) and best-arm identification (BAI).
Threshold Identification. We set threshold $\xi=0.5$. We use the sampling algorithm HDoC proposed by Kano et al.,, 2019: at time $t$, HDoC sample arm $a_{t}^{\text{HDoC}}$, defined as $a_{t}^{\text{HDoC}}=\operatornamewithlimits{arg\,max}_{a\in[K]}\hat{\mu}_{t}(a)+\sqrt{\frac{\log(t)}{2N_{t}(a)}}.$ We report the first time in which we have rejected $i$ regions as $\tau_{i}$, i.e. $\tau_{i}$ corresponds to the first time that we have labeled $i$ arms in Table B.2, and the final stopping time (i.e. $\tau_{4}$) in Table 1 in the main body of the text.
B.3 Runtime Testing.
@@
B.4 Additional Simulations with Nonstandard Distributions
To emphasize that PEAK performs well across all arbitrary data distributions over $[0,1]$, we provide additional results that test nonstandard distributions beyond exponential families. In particular, we test the following distributions, and provide additional results that demonstrate the robustness of PEAK’s performance.
  • Mixture Distributions for a Single Stream. In the single-stream case (Section 5.1), we test the mixture distribution, where observations are generated from $\text{Unif}(0,1)$, $\text{Beta}(1,1)$, and $\text{Bern}(0.5)$ with probability 1/3 each. Figure B.1 plots the width of the confidence sequence at time $t$ (i.e. the volume of the null hypothesis set we cannot reject at time $t$) formed by inverting PEAK, which empirically demonstrates that PEAK does not suffer (and is relatively better) than the other methods we test.
    Figure B.1: Confidence Sequence Widths in the Single Arm Setting for Mixture Distribution Arm
  • Mixture Distribution for BAI. For the multi-stream best-arm identification problem, we test distributions with adversarial heavy tails (outliers) in the best-arm identification setting by modifying our Beta distribution set-up. In our new simulations, we contaminate the distribution of arm 3 (second-best arm) with a point mass of $0.05$ at 1, and contaminate the distribution of arm 4 (best arm) with a point mass of $0.05$ at 0. To maintain the same $\mu$ across simulations, we choose Beta distributions with parameters $\alpha=1$ for both arms 3 and 4, and $\beta=0.43/0.52$, $\beta=0.24/0.71$ for arms 3 and 4 respectively. The updated Table provides empirical results regarding the stopping times, which align closely to the uncontaminated low-variance (uncontaminated Beta) setting.
    Type Stopping Rule Stopping Time
    PEAK 708.52 $\pm$ 266.34
    Beta Base 4686.1 $\pm$ 565.24
    Sub-B 3175.3 $\pm$ 905.70
    Hedged 500.12 $\pm$ 214.08
    PEAK 1318.14 $\pm$ 489.29
    Bern Base 4631.66 $\pm$ 896.51
    Sub-B 4322.36 $\pm$ 3638.25
    Hedged 1734.64 $\pm$ 858.66
    PEAK 705.72 $\pm$ 284.76
    Beta-Contaminated Base 4680.92 $\pm$ 676.86
    Sub-B 3376.22 $\pm$ 941.16
    Hedged 542.42 $\pm$ 246.24
    Table B.1: Stopping Times for BAI, including contaminated distribution example.
B.5 Additional Threshold Identification Results
We now provide the average stopping time for identifying each arm sequentially, where $\tau_{i}$ is the first time in which $i$ arms are been labeled as above/below the threshold.
Arm Type Stopping Rule $\tau_{1}$ $\tau_{2}$ $\tau_{3}$ $\tau_{4}$
Table B.2: Stopping Times $\tau$ for THR. Each $\tau_{i}$ represents the first time in which we have classified $i$ arms as above or below $\xi=0.5$. We report averaged stopping times and one standard error over 100 simulated sample paths. Bolded values represent the best and second-best average run-time respectively. Across all simulations, each stopping time results in the correct conclusion (i.e. rejects the incorrect hypothesis).
B.6 Visualization for Composite Hypothesis Testing for THR and BAI
We visualize Examples 1 and 2 in Section 4, providing simple intuition for the $W=2$ case. Below, we use $m_{t}^{*}$ to denote the global minimizer of $E_{t}(m)$, and $m_{t}^{i}$ to denote the minimizer of $E_{t}(m)$ for region $R_{i}$. These simple plots demonstrate the intuitive solutions implied by solving the KKT system of equations.
Figure B.2: Visualization of Minima within each region at time $t$, obtained by projecting the global minimizer of $E_{t}(m)$ onto the regions implied by THR (left) and BAI (right) for $W=2$ case. In both plots, the current global minima at time $t$ is contained in Region 1, and projected to obtain the minima in all other regions.